Twice Shy Industries has a debt-equity ratio of 1.8. Its WACC is 9.1 percent, and its cost of debt is 7.1 percent. The corporate tax rate is 35 percent a. What is the companys cost of equity capital? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Cost of equity capital 17.17% b. What is the companys unlevered cost of equity capital? (Do not round intermediate calculations Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Unlevered cost of equity capital c-1 What would the cost of equity be if the debt-equity ratio were 2? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Cost of equity c-2 What would the cost of equity be if the debt-equity ratio were 1.0? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Cost of equity c-3 What would the cost of equity be if the debt-equity ratio were zero? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Cost of equity
Answer #1

a)

WACC = (E/V) Requity + (D/V) Rdebt (1 – Tax )

Debt to equity = 1.8

lets assume debt = 1.8 and equity is 1

total V = 2.8

E/V = 1/2.8

D/V = 1.8/2.8

WACC = 0.091 = (1/2.8) * Requity + (1.8/2.8) * 0.071 * (1-0.35)

0.061332 = (1/2.8) * Requity

Requity = 17.17

b)

Requity = Runlevered + ( Runlevered – Rdebt )( D/E )(1 – Tax )

0.1717 = Runlevered + ( Runlevered - 0.071) * 1.8 * 0.65

Runlevered   = 11.74%

c-1)

We can use the same formula as mentioned above in b

Requity = Runlevered + ( Runlevered – Rdebt )( D/E )(1 – Tax )

= 0.1174 + (0.1174 - 0.071) * 2 * 0.65

= 17.77%

c-2)

At D/E = 1

Requity = Runlevered + ( Runlevered – Rdebt )( D/E )(1 – Tax )

= 0.1174 + (0.1174 - 0.071) * 1 * 0.65

= 14.75

c-3)

At D/E = 0

Requity = Runlevered + ( Runlevered – Rdebt )( D/E )(1 – Tax )

= 0.1174 + (0.1174 - 0.071) * 0 * 0.65

= 11.74%

Here Requity = Runlevered